I've done a few things to start waking up the IFO after it's week of conference-vacation.
PMC trans was at 0.679, aligned the input to the PMC, now it's up at 0.786.
MC transmission was very low, mostly from low PMC transmission. Anyhow, MC locked, WFS relieved so that it will re-acquire faster.
Many of the optics had drifted away. AS port had no fringing, and almost every optic was far away from it's driftmon set val. While putting the optics back to their driftmon spots, I noticed that some of the cds.servos had incorrect gain. Previously, I had just been using the ETMX servo, which had the correct gain, but the ITMs needed smaller gain, and some of the optics needed the gain to be negative rather than positive. So, now the script ..../scripts/SUS/DRIFT_MON/MoveOpticToMatchDriftMon.py has individually defined gains for the cds.servo.
Next up (after lunch) will be locking an aligning the arms. I still don't have MICH fringing at the AS port, so I suspect that the ASS will move some of the optics somewhat significantly (perhaps the input tip tilts, which I don't have DRIFT_MON for?) |